OpenStudy (anonymous):

-4(-3x-4)=100 solve for x....i have more than just this one to do. need "help"

pooja195 (pooja195):

Distribute whats -4 times -3?

OpenStudy (anonymous):

12

pooja195 (pooja195):

-4 times -4?

OpenStudy (anonymous):

16

pooja195 (pooja195):

so now you have 12x+16=100 now subtract 16 from both sides

pooja195 (pooja195):

whats 100 - 16?

OpenStudy (anonymous):

so 100-16?

pooja195 (pooja195):

lol yes :)

OpenStudy (anonymous):

so 84

pooja195 (pooja195):

yes now 84/12 is?

OpenStudy (anonymous):

7

pooja195 (pooja195):

so x=?

OpenStudy (anonymous):

7

pooja195 (pooja195):

well done :)

OpenStudy (anonymous):

oh okay i got it now...haha thanks...lol i kinda fell asleep in class when my math teacher was going over it so yeah...lol

pooja195 (pooja195):

LOL i know how you feel im not a big fan of math either :P but your welcome :)

OpenStudy (anonymous):

-4(7-2x)+4=-88

OpenStudy (thatgingeronfire):

-4*7=? -4*-2=?x

OpenStudy (anonymous):

its cool i figured it out